what is the expected volume of 10 grams of oxygen gas at STP? given that the volume of 1 mole of a gas at STP is 22.4 Liters
1
Expert's answer
2020-10-28T08:40:26-0400

M (O2)=32 /mol

n=m/M

n (O2)=10/32=0.3 mol

V (O2) = 0.3 x 22.4 = 6.72 l